OpenStudy (anonymous):

a golf ball is packed into a cubical box so that it just fits. what percentage of the box is unfilled?

OpenStudy (gorv):

volume of box - volume of ball

OpenStudy (gorv):

divide by volume of box

OpenStudy (anonymous):

That is what I did. Isn't the 2x radius the same as length, width, height of the cubic box?

OpenStudy (gorv):

foe a cube length=breadth=height=l volume=l^3 volume of ball=4/3*pi*r^3

OpenStudy (anonymous):

My answer is 476.4%

OpenStudy (anonymous):

Is that how we do it?

OpenStudy (gorv):

(l^3-4/3*pi*r^3)/l^3

OpenStudy (gorv):

ans will be less than 100%

OpenStudy (anonymous):

You'll get a different answer (0.869)

OpenStudy (gorv):

numerator will be less then denominator let ans is 0.99999 for % calculation 0.99999*100=99.999 so it can not go above 100 give me the value i will solve fr u

OpenStudy (anonymous):

They haven't given a value but one can conclude that if the radius is 5 then the height, length and width are all equal to 10 as the golf ball perfectly fits the cubic box

OpenStudy (anonymous):

You would need to draw it

OpenStudy (anonymous):

 to find the percentage of the box that is unfilled you can just find the percentage of the box that is filled and subtract that from 100%

OpenStudy (anonymous):

to find this, you can assume the radius of the sphere to be 1, which makes the side length of the box 2, your reasoning above is correct.

OpenStudy (anonymous):

area of a cube = side^3 = 8 area of a sphere = 4/3 pi*r^3 = 4/3 pi

OpenStudy (anonymous):

so the percent filled is \[\frac {\frac{ 4 }{ 3 }\times \pi}{8}\]

OpenStudy (anonymous):

and from there you can find the percent empty @Chad123 you see how this works?

OpenStudy (gorv):

@Peter14 we have to calculate empty space ratio so (8-4*pi/3)/8

OpenStudy (anonymous):

Thanks guys

OpenStudy (anonymous):

yes, @gorv I specified earlier that you subtract the percent filled space from 100

OpenStudy (anonymous):

So we don't have to convert the answer into a %?

OpenStudy (gorv):

@Peter14 actually i didn't see that sorry for that